# Constrained Online Convex Optimization with Polyak Feasibility Steps

### v1

@@ 1.1.3 Constrained Optimization
Our approach is inspired by a line of literature in constrained (offline) optimization that uses the Polyak step-size to ensure convergence to the feasible set, e.g. Polyak (2001); Nedić (2011); Nedić and Necoara (2019); Necoara and Singh (2022). However, we point out two key difficulties that arise in the OCO setting: (a) the suboptimality gap of the iterates $f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{x}^{\star})$ is not guaranteed to be non-negative, and (b) the constraint feedback is at the played action $\mathbf{x}_{t}$ and not at the “intermediate” iterate (labeled $\mathbf{y}_{t}$ in Algorithm 1). Challenge (a) is particularly difficult to handle because the analysis approach used by Nedić (2011) (and following works) relies on the suboptimality gap of the iterates being non-negative. As a result, we require a new analysis approach. However, our approach does suffer a larger dependence on the problem parameters (such as the subgradient bound $G_{g}$), which can be seen as the “cost” of the adversarial online setting. To handle challenge (b), we use the first-order approximation of the cost function $g_{t}+\mathbf{s}_{t}^{\top}(\mathbf{y}_{t}-\mathbf{x}_{t})$ in the Polyak step-size instead of the true cost function $g(\mathbf{y}_{t})$. This first-order approximation avoids the need for constraint information at the intermediate iterate $\mathbf{y}_{t}$, while maintaining the advantageous properties of the Polyak step-size (see Section 3.4 for the details). Overall, we believe that these analysis techniques will be of interest to the constrained optimization literature, beyond just OCO.
@@ 2 Problem Setup
We will use the standard assumptions that the cost functions have bounded gradients and that the action set is bounded. These are stated precisely in the following.
@@
There exists positive real $G_{f}$ such that $\|\nabla f_{t}(\mathbf{x})\|\leq G_{f}$ for all $\mathbf{x}\in R\mathbb{B}$ and $t\in[T]$.
@@
Following Mahdavi et al. (2012), we study the setting where the action set is defined by a functional inequality constraint $\mathcal{X}=\{\mathbf{x}\in\mathbb{R}^{d}:g(\mathbf{x})\leq 0\}$, with $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$ being non-smooth and convex. In the following, we assume that the constraint function has bounded subgradients (Assumption 3) and that the subgradient norm is lower-bounded near the boundary (Assumption 4). These assumptions are also used by Mahdavi et al. (2012) and Jenatton et al. (2016). Since the constraint function is non-smooth, this setting can be extended to multiple constraints $g_{1},..,g_{m}$ by taking $g(\mathbf{x})=\max_{i\in[m]}g_{i}(\mathbf{x})$.
@@
There exists positive reals $\sigma,\epsilon$ such that $\|\partial g(\mathbf{x})\|\geq\sigma$ for all $\mathbf{x}\in\mathbb{R}^{d}$ where $g(\mathbf{x})=-\epsilon$.
@@ 3 Algorithm
Algorithm 1 OGD with Polyak Feasibility Steps
0:  initial action $\mathbf{x}_{1}\in\mathbb{R}^{d}$, step size $\eta\in\mathbb{R}$, tightening $\rho\in\mathbb{R}$.
1:  for $t=1,2,...,T$ do
2:     Play $\mathbf{x}_{t}$ and receive $f_{t}$.
5:     Polyak feasibility step: [^4] The projection on to the ball can be implemented by scaling the point according to its norm. Specifically, for $\mathbf{x}\in\mathbb{R}^{d}$, it holds that $\Pi_{R\mathbb{B}}(\mathbf{x})=\min\left(1,\frac{R}{\|\mathbf{x}\|}\right)\mathbf{x}$.$\mathbf{x}_{t+1}=\Pi_{R\mathbb{B}}\left(\mathbf{y}_{t}-\frac{[g_{t}+\mathbf{s}_{t}^{\top}(\mathbf{y}_{t}-\mathbf{x}_{t})+\rho]_{+}}{\|\mathbf{s}_{t}\|^{2}}\mathbf{s}_{t}\right)$.
6:  end for
@@ 3.1 Description
The design of our Polyak feasibility step is motivated by the classical Polyak step-size Polyak (1969). In unconstrained convex optimization, using subgradient descent with the Polyak step-size is known to be optimal Boyd et al. (2003), and therefore it is a natural choice to ensure strong feasibility guarantees when applied to the constraint function. This classical step-size uses the function value and subgradient at the current iterate to approximate the optimal step-size in each update. In our setting, this would require the constraint function value at the “intermediate iterate” $\mathbf{y}_{t}$, which is not available. We address this next.
@@
Another difficulty that arises in our setting is that the gradient descent step in line 4 might push the sequences of actions out of the feasible set. Indeed, the cost functions are chosen adversarially and therefore we have no guarantees about the direction of the gradient in each round. Therefore, to ensure that the actions are feasible, we use a tightened version of the constraint function $g(\mathbf{x})+\rho$ where $\rho>0$ is a tightening parameter that is to be chosen appropriately. Given that the constraint function is Lipschitz (via Assumption 3), this ensures that there is a “buffer zone” between the points that satisfy the tightened constraint $g(\mathbf{x})+\rho\leq 0$ and the boundary of the true feasible set defined by $g(\mathbf{x})\leq 0$. Therefore, by choosing $\rho$ proportional to the cost step-size $\eta$, we can ensure that the actions are feasible despite the adversarially-chosen cost gradients. Note that tightening the constraint in this manner is a common technique in the constrained OCO literature, e.g. Mahdavi et al. (2012); Jenatton et al. (2016).
@@ 3.2 Guarantees
The following theorem shows that, with appropriate choice of algorithm parameters, Algorithm 1 enjoys $\mathcal{O}(\sqrt{T})$ regret in terms of the horizon $T$, and anytime constraint satisfaction.
Theorem 1 (Main Result).
Let Assumptions 1, 2, 3 and 4 hold. Then, playing Algorithm 1 with $g(\mathbf{x}_{1})\leq-\rho$,
$\eta=\frac{\xi\epsilon}{G_{f}G_{g}\sqrt{T}}\qquad\text{and}\qquad\rho=\frac{\epsilon}{\sqrt{T}},$
ensures that,
$\mathrm{Reg}_{T}$ $\leq\left(\frac{G_{f}G_{g}R^{2}}{2\xi\epsilon}+\frac{G_{f}\xi\epsilon}{2G_{g}}+\frac{G_{f}\epsilon}{\sigma}\right)\sqrt{T},$
@@
where, $\xi=1-\sqrt{1-\frac{\sigma^{2}}{G_{g}^{2}}}$.
One drawback of the guarantee in Theorem 1 is that it requires the knowledge of a strictly feasible point as the initial action, i.e. $g(\mathbf{x}_{1})\leq-\rho$. To address this, we give Corollary 1 below that only requires $\mathbf{x}_{1}\in R\mathbb{B}$ and ensures that $g(\mathbf{x}_{t})\leq 0$ for all $t\geq t_{0}$ where $t_{0}=\mathcal{O}(\log(T))$. Note that $\mathbf{x}_{1}\in R\mathbb{B}$ can always be satisfied given an arbitrary $\mathbf{x}\in\mathbb{R}^{d}$, by taking $\mathbf{x}_{1}=\min\left(\frac{R}{\|\mathbf{x}\|},1\right)\mathbf{x}$. Corollary 1 also shows that when $T$ is sufficiently large, then the algorithm ensures that there is cumulative constraint satisfaction $\sum_{t=1}^{T}g(\mathbf{x}_{t})\leq 0$. Note that the requirement that $T$ is sufficiently large is also used by prior work that shows cumulative constraint satisfaction Mahdavi et al. (2012); Jenatton et al. (2016).
@@
Assume the same as Theorem 1. Consider Algorithm 1 with $\mathbf{x}_{1}\in R\mathbb{B}$, $\eta=\frac{\xi\epsilon}{2G_{f}G_{g}\sqrt{T}}$ and $\rho=\frac{\epsilon}{\sqrt{T}}$. Then, it holds that,
@@
$g(\mathbf{x}_{t})$ $\leq 0\quad\forall t\geq 1+\frac{2G_{g}^{2}}{\sigma^{2}}\log\left(\frac{2G_{g}R\sqrt{T}}{\epsilon}\right).$
Furthermore, when $\sqrt{T}\geq\frac{2RG_{g}}{\epsilon\xi}$ it additionally holds that,
@@
Lastly, we give another corollary below (Corollary 2) that allows a small amount of violation in each round, while eliminating the dependence on $G_{g}$ in the regret bound. In this case, the regret bound is $RG_{f}\sqrt{T}$, which matches what is attained by online gradient descent using a full projection on to the action set in each round Zinkevich (2003). At the same time, the constraint violation satisfies $g(x_{t})=\mathcal{O}(\frac{1}{\sqrt{T}})$.
@@
Assume the same as in Theorem 1. Consider Algorithm 1 with $\mathbf{x}_{1}\in R\mathbb{B}$, $\eta=\frac{G_{f}}{R\sqrt{T}}$ and $\rho=0$. Then, it holds that,
$\mathrm{Reg}_{T}$ $\leq RG_{f}\sqrt{T},$
$g(\mathbf{x}_{t})$ $\leq\exp\left(-\frac{\sigma^{2}}{2G_{g}^{2}}t\right)RG_{g}+\frac{G^{2}_{f}G_{g}}{\xi R\sqrt{T}}\quad\forall t\in[T]$
@@ 3.3 Regret Analysis
We start with Term I. The key observation is that the Polyak feasibility step (line 5) can be equivalently defined as the projection on to the halfspace $\mathcal{H}_{t}$ in (5) (with the additional projection on to the ball). In fact, this halfspace $\mathcal{H}_{t}$ contains the tightened feasible set $\mathcal{X}_{\rho}$ and therefore projecting on to it will not increase the distance to any point in $\mathcal{X}_{\rho}$. This is stated precisely in Fact 1.
@@
The update for $\mathbf{x}_{t+1}$ in line 5 of Algorithm 1 is equivalent to $\mathbf{x}_{t+1}=\Pi_{R\mathbb{B}}(\Pi_{\mathcal{H}_{t}}(\mathbf{y}_{t+1}))$ where,
@@
$\mathrm{Term\ I}\leq\frac{R^{2}}{2\eta}+\frac{\eta}{2}G_{f}^{2}T.$ (7)
We defer the complete proof of (7) to Appendix B.2.
Now, we look at Term II. We use Lemma 1 below, which provides a bound on the distance to $\mathcal{X}_{\rho}$ in terms of $g$. This lemma is conceptually similar to Theorem 7 in Mahdavi et al. (2012). However, it is more general in the sense that it is a general error bound on the constraint function, whereas Theorem 7 in Mahdavi et al. (2012) only provides a bound on the difference in costs between a point in the tightened feasible set and the original feasible set. The proof of Lemma 1 is given in Appendix B.1.
@@
Combining the bounds on Term I and Term II and using the choice of $\eta,\alpha$ give the complete regret bound in Theorem 1.
@@ 3.4 Feasibility Analysis
Let Assumption 3 and 4 hold, and suppose that $\rho\in[0,\epsilon]$. Furthermore, consider the $\rho$-sublevel set of $g$, $\mathcal{X}_{\rho}:=\{\mathbf{x}\in\mathbb{R}^{d}:g(\mathbf{x})+\rho\leq 0\}$. Consider any $\mathbf{x}\in R\mathbb{B}$ and $\mathbf{s}\in\partial g(\mathbf{x})$, and let,
@@
where (a) uses the Pythagorean theorem of the projection, (b) uses that $\mathbf{s}^{\top}(\mathbf{x}-\mathbf{v})\geq g_{\rho}(\mathbf{x})-g_{\rho}(\mathbf{v})\geq g_{\rho}(\mathbf{x})$ due to convexity of $g_{\rho}$ and the fact that $\mathbf{v}\in\mathcal{X}_{\rho}$, (c) uses that $\|\mathbf{s}\|\leq G_{g}$ due to Assumption 3, and (d) uses Lemma 1. ∎
Note that Lemma 2 uses the “exact” Polyak step-size in the sense that the numerator of (9) uses the exact constraint function value at the previous iterate $g(\mathbf{x})$. This differs from the update in line 5 in Algorithm 1, which uses the first-order approximation of the function value at $\mathbf{y}_{t}$ because the algorithm does not have access to the exact function value at $\mathbf{y}_{t}$. In order to handle this disparity, we introduce the “fictitious” iterate $\mathbf{z}_{t+1}$,
@@
Importantly, the update for $\mathbf{z}_{t+1}$ matches the form in (9) and therefore we can directly apply Lemma 2 analyze $\mathbf{z}_{t+1}$.
@@
We start with Term I. Using the “halfspace representation” of the Polyak feasibility step (as in Fact 1), we can write $\mathbf{x}_{t+1}=\Pi_{R\mathbb{B}}(\Pi_{\mathcal{H}_{t}}(\mathbf{y}_{t}))$ and $\mathbf{z}_{t+1}=\Pi_{R\mathbb{B}}(\Pi_{\mathcal{H}_{t}}(\mathbf{x}_{t}))$ where $\mathcal{H}_{t}$ is defined in (5). Therefore, it follows that,
$\mathrm{Term\ I}$ $=\|\Pi_{R\mathbb{B}}(\Pi_{\mathcal{H}_{t}}(\mathbf{y}_{t}))-\Pi_{R\mathbb{B}}(\Pi_{\mathcal{H}_{t}}(\mathbf{x}_{t}))\|$
$\leq\|\Pi_{\mathcal{H}_{t}}(\mathbf{y}_{t})-\Pi_{\mathcal{H}_{t}}(\mathbf{x}_{t})\|$
$\leq\|\mathbf{y}_{t}-\mathbf{x}_{t}\|$
$\leq\eta G_{f},$
@@
Next, we look at Term II. Applying Lemma 2 yields,
$(\mathrm{Term\ II})^{2}\leq\gamma\ \mathrm{dist}^{2}(\mathbf{x}_{t},\mathcal{X}_{\rho}),$
where $\gamma=1-\frac{\sigma^{2}}{G_{g}^{2}}$.
@@
$\leq\gamma^{t/2}\mathrm{dist}(\mathbf{x}_{1},\mathcal{X}_{\rho})+\frac{\eta G_{f}}{1-\sqrt{\gamma}},$
where we apply the bound recursively in the second line. Then, applying the the subgradient bound (Assumption 3),
$\begin{split}g(\mathbf{x}_{t})&=g(\mathbf{x}_{t})-g(\Pi_{\mathcal{X}_{\rho}}(\mathbf{x}_{t}))+g(\Pi_{\mathcal{X}_{\rho}}(\mathbf{x}_{t}))\\ &\leq G_{g}\mathrm{dist}(\mathbf{x}_{t},\mathcal{X}_{\rho})-\rho\\ &\leq G_{g}\gamma^{(t-1)/2}\mathrm{dist}(\mathbf{x}_{1},\mathcal{X}_{\rho})+\frac{\eta G_{g}G_{f}}{1-\sqrt{\gamma}}-\rho,\end{split}$ (11)
The guarantees in Theorem 1 follows from (11) by applying the choice of $\mathbf{x}_{1}$, $\eta$ and $\rho$, as this ensures that $\mathrm{dist}(\mathbf{x}_{1},\mathcal{X}_{\rho})=0$ and that $\frac{\eta G_{g}G_{f}}{1-\sqrt{\gamma}}=\rho$. Also, the guarantees in Corollary 1 follow by applying the bound $\mathrm{dist}(\mathbf{x}_{1},\mathcal{X}_{\rho})\leq R$ and then finding when the right-hand side is less than 00 (with the specified choice of $\eta$ and $\rho$). Lastly, the guarantees in Corollary 2 follow by simply taking $\rho=0$.

### v2

@@ 1.1.3 Constrained Optimization
Our approach is inspired by a line of literature in constrained (offline) optimization that uses the Polyak step-size to ensure convergence to the feasible set, e.g. (Polyak, 2001; Nedić, 2011; Nedić & Necoara, 2019; Necoara & Singh, 2022). However, we point out two key difficulties that arise in the OCO setting: (a) the suboptimality gap of the iterates $f_{t}(\mathbf{x}_{t})-f_{t}(\mathbf{x}^{\star})$ is not guaranteed to be non-negative, and (b) the constraint feedback is at the played action $\mathbf{x}_{t}$ and not at the “intermediate” iterate (labeled $\mathbf{y}_{t}$ in Algorithm 1). Challenge (a) is particularly difficult to handle because the analysis approach used by Nedić (2011) (and following works) relies on the suboptimality gap of the iterates being non-negative. As a result, we require a new analysis approach. However, our approach does suffer a larger dependence on the problem parameters (such as the subgradient bound $G_{g}$), which can be seen as the “cost” of the adversarial online setting. To handle challenge (b), we use the first-order approximation of the cost function $g_{t}+\mathbf{s}_{t}^{\top}(\mathbf{y}_{t}-\mathbf{x}_{t})$ in the Polyak step-size instead of the true cost function $g(\mathbf{y}_{t})$. This first-order approximation avoids the need for constraint information at the intermediate iterate $\mathbf{y}_{t}$, while maintaining the advantageous properties of the Polyak step-size (see Section 3.4 for the details). Such a first-order approximation is also advantageous in the offline setup as seen in the recent work of Singh & Necoara (2025), but the analysis in Singh & Necoara (2025) does not extend to the OCO setting because it requires that the suboptimality gap is non-negative (i.e. $f(x_{t})-f(x^{\star})\geq 0$) and that the cost functions have Lipschitz gradients. Related algorithm designs have also been considered in the literature on constrained variational inequalities (Zhang et al., 2025), although under the assumption that the constraint functions have Lipschitz gradients (which does not hold in our setting). We further discuss the differences between our algorithm and the algorithm in Zhang et al. (2025) in Appendix B.
@@ 2 Problem Setup
We will use the standard assumptions that the action set is bounded and that the cost functions have bounded gradients. These are stated precisely in the following.
@@
There exists positive real $G_{f}$ such that $\|\nabla f_{t}(\mathbf{x})\|\leq G_{f}$ for all $\mathbf{x}\in R\mathbb{B}$ and $t\in[T]$. [^4] This assumption can be weakened slightly for our guarantees of anytime constraint satisfaction (i.e. Corollary 1). Specifically, we only need that $\|\nabla f_{t}(\mathbf{x})\|\leq G_{f}$ for all $\mathbf{x}\in\mathcal{X}$ (rather than all $\mathbf{x}\in R\mathbb{B}$). We discuss this later in Remark 2.
@@
Following Mahdavi et al. (2012), we study the setting where the action set is defined by a functional inequality constraint $\mathcal{X}=\{\mathbf{x}\in\mathbb{R}^{d}:g(\mathbf{x})\leq 0\}$, with $g:\mathbb{R}^{d}\rightarrow\mathbb{R}$ being non-smooth and convex. In the following, we assume that the constraint function has bounded subgradients (Assumption 3) and that the subgradient norm is lower-bounded near the boundary (Assumption 4). These assumptions are also used by Mahdavi et al. (2012) and Jenatton et al. (2016). We also note in Remark 1 that Assumption 4 is implied by Slater’s condition (i.e. the existence of a strictly-feasible point) provided that the other assumptions hold. Lastly, note that since the constraint function is non-smooth, this setting can be extended to multiple constraints $g_{1},..,g_{m}$ by taking $g(\mathbf{x})=\max_{i\in[m]}g_{i}(\mathbf{x})$.
@@
There exists positive reals $\sigma,\epsilon$ such that $\mathcal{X}^{\prime}=\{\mathbf{x}\in\mathbb{R}^{d}:g(\mathbf{x})=-\epsilon\}$ is nonempty and $\|\partial g(\mathbf{x})\|\geq\sigma$ for all $\mathbf{x}\in\mathcal{X}^{\prime}$.
Remark 1.
Assumption 4 holds if Assumption 1 holds and there exists a strictly-feasible point (i.e. Slater’s condition). We show this in Appendix C.1.
@@ 3.1 Description
The design of our Polyak feasibility step is motivated by the classical Polyak step-size (Polyak, 1969). In unconstrained convex optimization, using subgradient descent with the Polyak step-size is known to be optimal (Boyd et al., 2003), and therefore it is a natural choice to ensure strong feasibility guarantees when applied to the constraint function. This classical step-size uses the function value and subgradient at the current iterate to approximate the optimal step-size in each update. In our setting, this would require the constraint function value at the “intermediate iterate” $\mathbf{y}_{t}$, which is not available. We address this next.
@@
Another difficulty that arises in our setting is that the gradient descent step in line 4 might push the sequences of actions out of the feasible set. Indeed, the cost functions are chosen adversarially and therefore we have no guarantees about the direction of the gradient in each round. Therefore, to ensure that the actions are feasible, we use a tightened version of the constraint function $g(\mathbf{x})+\rho$ where $\rho>0$ is a tightening parameter that is to be chosen appropriately. Given that the constraint function is Lipschitz (via Assumption 3), this ensures that there is a “buffer zone” between the points that satisfy the tightened constraint $g(\mathbf{x})+\rho\leq 0$ and the boundary of the true feasible set defined by $g(\mathbf{x})\leq 0$. Therefore, by choosing $\rho$ proportional to the cost step-size $\eta$, we can ensure that the actions are feasible despite the adversarially-chosen cost gradients. Note that tightening the constraint in this manner is a common technique in the constrained OCO literature, e.g. (Mahdavi et al., 2012; Jenatton et al., 2016).
@@
Algorithm 1 OGD with Polyak Feasibility Steps
0:  initial action $\mathbf{x}_{1}\in\mathbb{R}^{d}$, step size $\eta>0$, tightening $\rho\geq 0$.
1:  for $t=1,2,...,T$ do
2:     Play $\mathbf{x}_{t}$ and receive $f_{t}$.
5:     Polyak feasibility step: [^5] When $\mathbf{s}_{t}=\mathbf{0}$, we take the update to be $\mathbf{x}_{t+1}=\Pi_{R\mathbb{B}}\left(\mathbf{y}_{t}\right)$.$\mathbf{x}_{t+1}=\Pi_{R\mathbb{B}}\left(\mathbf{y}_{t}-\frac{[g_{t}+\mathbf{s}_{t}^{\top}(\mathbf{y}_{t}-\mathbf{x}_{t})+\rho]_{+}}{\|\mathbf{s}_{t}\|^{2}}\mathbf{s}_{t}\right)$.
6:  end for
@@ 3.2 Guarantees
In this section, we give the regret guarantees and constraint satisfaction guarantees for Algorithm 1. In particular, the following theorem (Theorem 1) gives a bound on the regret and constraint violation for an arbitrary choice of algorithm parameters $\eta,\rho$. We will then show how these algorithm parameters can be chosen to get several different guarantees.
Theorem 1.
Let Assumptions 1, 2, 3 and 4 hold. Then, playing Algorithm 1 with $\eta>0$ and $\rho\in[0,\epsilon]$ ensures that,
$\mathrm{Reg}_{T}\leq\frac{2R^{2}}{\eta}+\frac{\eta}{2}G_{f}^{2}T+\frac{G_{f}\rho}{\sigma}T,$
$g(\mathbf{x}_{t})\leq G_{g}\gamma^{(t-1)/2}\mathrm{dist}(\mathbf{x}_{1},\mathcal{X}_{\rho})+\frac{\eta G_{g}G_{f}}{1-\sqrt{\gamma}}-\rho,\ \forall t$
where $\mathcal{X}_{\rho}=\{x\in\mathbb{R}^{d}:g(\mathbf{x})\leq-\rho\}$ and $\gamma=1-\frac{\sigma^{2}}{G_{g}^{2}}$.
In the following, Corollary 1 specifies a choice of $\eta$ and $\rho$ that ensures $\mathcal{O}(\sqrt{T})$ regret and anytime constraint satisfaction. Note that this result requires that the initial action $\mathbf{x}_{1}$ is strictly-feasible. We also point out that these guarantees still hold under a slightly weaker set of assumptions, which we discuss in Remark 2.
Corollary 1.
Let Assumptions 1, 2, 3 and 4 hold. Suppose that $g(\mathbf{x}_{1})\leq-\alpha$ for some $\alpha\in(0,\epsilon]$, and let $\rho=\frac{\alpha}{\sqrt{T}}$ and $\eta=\frac{\xi\rho}{G_{f}G_{g}}$, where $\xi=1-\sqrt{\gamma}$. It follows that the actions chosen by Algorithm 1 satisfy,
$\mathrm{Reg}_{T}$ $\leq\left(\frac{2G_{f}G_{g}R^{2}}{\xi\alpha}+\frac{G_{f}\xi\alpha}{2G_{g}}+\frac{G_{f}\alpha}{\sigma}\right)\sqrt{T},$
@@
Remark 2.
The guarantees in Corollary 1 also hold under a slightly weaker set of assumptions. In particular, we only need that $\|\nabla f_{t}(\mathbf{x})\|\leq G_{f}$ for all $\mathbf{x}\in\mathcal{X}$ (rather than all $\mathbf{x}\in R\mathbb{B}$). We show this in Appendix C.4.
Next, we give Corollary 2, which shows that, when $\mathbf{x}_{1}$ is in $R\mathbb{B}$ (and therefore is not necessarily feasible), then the algorithm can ensure anytime constraint satisfaction after $\mathcal{O}(\log(T))$ rounds. Corollary 2 also shows that when $T$ is sufficiently large, then the algorithm ensures that there is cumulative constraint satisfaction $\sum_{t=1}^{T}g(\mathbf{x}_{t})\leq 0$. Note that the requirement that $T$ is sufficiently large is also used by prior work that shows cumulative constraint satisfaction (Mahdavi et al., 2012; Jenatton et al., 2016).
@@
Let Assumptions 1, 2, 3 and 4 hold. Consider Algorithm 1 with $\mathbf{x}_{1}\in R\mathbb{B}$, $\eta=\frac{\xi\epsilon}{2G_{f}G_{g}\sqrt{T}}$ and $\rho=\frac{\epsilon}{\sqrt{T}}$. Then, it holds that,
@@
$g(\mathbf{x}_{t})$ $\leq 0\quad\forall t\geq 1+\frac{2G_{g}^{2}}{\sigma^{2}}\log\left(\frac{4G_{g}R\sqrt{T}}{\epsilon}\right).$
Furthermore, when $\sqrt{T}\geq\frac{4RG_{g}}{\epsilon\xi}$ it additionally holds that,
@@
Lastly, we give another corollary below (Corollary 3) that allows a small amount of violation in each round, while eliminating the dependence on $G_{g}$ in the regret bound. In this case, the regret bound is $2RG_{f}\sqrt{T}$, which matches what is attained by online gradient descent using a full projection on to the action set in each round. At the same time, the constraint violation satisfies $g(x_{t})=\mathcal{O}(\frac{1}{\sqrt{T}})$.
@@
Let Assumptions 1, 2, 3 and 4 hold. Consider Algorithm 1 with $\mathbf{x}_{1}\in R\mathbb{B}$, $\eta=\frac{2R}{G_{f}\sqrt{T}}$ and $\rho=0$. Then, it holds that,
$\mathrm{Reg}_{T}$ $\leq 2RG_{f}\sqrt{T},$
$g(\mathbf{x}_{t})$ $\leq 2RG_{g}\exp\left(-\frac{\sigma^{2}(t-1)}{2G_{g}^{2}}\right)+\frac{2RG_{g}}{\xi\sqrt{T}}\quad\forall t\in[T]$
@@ 3.3 Regret Analysis
We start with Term I. The key observation is that the Polyak feasibility step (line 5) can be equivalently defined as the projection on to the halfspace $\mathcal{H}_{t}$ in (5) below (with the additional projection on to the ball). In fact, this halfspace $\mathcal{H}_{t}$ contains the tightened feasible set $\mathcal{X}_{\rho}$ and therefore projecting on to it will not increase the distance to any point in $\mathcal{X}_{\rho}$. This is stated precisely in Fact 1.
@@
The update for $\mathbf{x}_{t+1}$ in line 5 of Algorithm 1 is equivalent to $\mathbf{x}_{t+1}=\Pi_{R\mathbb{B}}(\Pi_{\mathcal{H}_{t}}(\mathbf{y}_{t}))$ where,
@@
$\mathrm{Term\ I}\leq\frac{2R^{2}}{\eta}+\frac{\eta}{2}G_{f}^{2}T.$ (7)
We defer the complete proof of (7) to Appendix C.3.
Now, we look at Term II. We use Lemma 1 below, which provides a bound on the distance to $\mathcal{X}_{\rho}$ in terms of $g$. This lemma is conceptually similar to Theorem 7 in Mahdavi et al. (2012). However, it is more general in the sense that it is a general error bound on the constraint function, whereas Theorem 7 in Mahdavi et al. (2012) only provides a bound on the difference in costs between a point in the tightened feasible set and the original feasible set. The proof of Lemma 1 is given in Appendix C.2.
@@
Combining the bounds on Term I and Term II yields,
$\mathrm{Reg}_{T}\leq\frac{2R^{2}}{\eta}+\frac{\eta}{2}G_{f}^{2}T+\frac{G_{f}\rho}{\sigma}T,$
which matches the bound in Theorem 1.
@@ 3.4 Feasibility Analysis
Let Assumption 3 and 4 hold, and suppose that $\rho\in[0,\epsilon]$. Furthermore, consider the $\rho$-sublevel set of $g$, $\mathcal{X}_{\rho}:=\{\mathbf{x}\in\mathbb{R}^{d}:g(\mathbf{x})+\rho\leq 0\}$. Consider any $\mathbf{x}\in R\mathbb{B}$ and $\mathbf{s}\in\partial g(\mathbf{x})$ such that $\mathbf{s}\neq\mathbf{0}$, and let,
@@
where (a) uses the Pythagorean theorem of the projection, (b) uses that $\mathbf{s}^{\top}(\mathbf{x}-\mathbf{v})\geq g_{\rho}(\mathbf{x})-g_{\rho}(\mathbf{v})\geq g_{\rho}(\mathbf{x})$ due to the fact that $\mathbf{s}\in\partial g(\mathbf{x})$ and $\mathbf{v}\in\mathcal{X}_{\rho}$, (c) uses that $\|\mathbf{s}\|\leq G_{g}$ due to Assumption 3, and (d) uses Lemma 1. ∎
Using Lemma 2, we will show that the distance between the action $\mathbf{x}_{t}$ and the tightened feasible set $\mathcal{X}_{\rho}$ will not increase too much in each round, and therefore that $\mathbf{x}_{t}$ always stays close to $\mathcal{X}_{\rho}$. The key difficulty in doing so is that Lemma 2 uses the “exact” Polyak step-size, while Algorithm 1 uses the first-order approximation of the constraint value in the Polyak step-size (i.e. as in (2)). In order to handle this disparity, we introduce the “fictitious” iterate $\mathbf{z}_{t+1}$. We define $\mathbf{z}_{t+1}$ such that $\mathbf{z}_{t+1}=\mathbf{x}_{t}$ when $\mathbf{s}_{t}=\mathbf{0}$, and otherwise take,
@@
We can interpret $\mathbf{z}_{t+1}$ as a feasibility step that is taken directly from the previous action $\mathbf{x}_{t}$. Importantly, the update for $\mathbf{z}_{t+1}$ matches the form in (9) and therefore we will be able to directly apply Lemma 2 to analyze $\mathbf{z}_{t+1}$.
@@
Figure 1: Experiment results for our algorithm (labeled PFS), alongside the algorithm from Yu et al. (2017) (labeled DPP). We also include a version of the algorithm in Yu et al. (2017) where the constraint is tightened (labeled DPP-T). The points indicate the average over 30 trials and the error bars and shading are ±$\pm 1$ standard deviation.
We start with Term I. First, note that $\mathbf{z}_{t+1}$ can be written as $\mathbf{z}_{t+1}=\Pi_{R\mathbb{B}}(\Pi_{\mathcal{H}_{t}}(\mathbf{x}_{t}))$, where $\mathcal{H}_{t}$ is the halfspace defined in (5). Therefore, it follows that,
$\begin{split}\mathrm{Term\ I}&=\|\Pi_{R\mathbb{B}}(\Pi_{\mathcal{H}_{t}}(\mathbf{y}_{t}))-\Pi_{R\mathbb{B}}(\Pi_{\mathcal{H}_{t}}(\mathbf{x}_{t}))\|\\ &\leq\|\Pi_{\mathcal{H}_{t}}(\mathbf{y}_{t})-\Pi_{\mathcal{H}_{t}}(\mathbf{x}_{t})\|\\ &\leq\|\mathbf{y}_{t}-\mathbf{x}_{t}\|\\ &=\eta\|\nabla f_{t}(\mathbf{x}_{t})\|\\ &\leq\eta G_{f},\end{split}$ (11)
@@
Next, we look at Term II. If $\mathbf{s}_{t}=\mathbf{0}$, then it holds that $\mathbf{z}_{t+1}=\mathbf{x}_{t}$ and $g(\mathbf{x}_{t})=\min_{\mathbf{x}}g(\mathbf{x})\leq-\rho$, which implies that $\mathrm{Term\ II}=0$. Alternatively, if $\mathbf{s}_{t}\neq\mathbf{0}$, then applying Lemma 2 yields $(\mathrm{Term\ II})^{2}\leq\gamma\ \mathrm{dist}^{2}(\mathbf{x}_{t},\mathcal{X}_{\rho})$, where $\gamma:=1-\frac{\sigma^{2}}{G_{g}^{2}}$. Therefore, it holds in any case that,
$\mathrm{Term\ II}\leq\sqrt{\gamma}\mathrm{dist}(\mathbf{x}_{t},\mathcal{X}_{\rho})$
@@
$\leq\gamma^{t/2}\mathrm{dist}(\mathbf{x}_{1},\mathcal{X}_{\rho})+\eta G_{f}\sum_{s=0}^{t-1}(\sqrt{\gamma})^{s}$
$\leq\gamma^{t/2}\mathrm{dist}(\mathbf{x}_{1},\mathcal{X}_{\rho})+\frac{\eta G_{f}}{1-\sqrt{\gamma}}$
where we apply the bound recursively in the second line, and use the fact that $\gamma\in[0,1)$ in the third line. Then, applying the the subgradient bound (Assumption 3),
$g(\mathbf{x}_{t})$ $=g(\mathbf{x}_{t})-g(\Pi_{\mathcal{X}_{\rho}}(\mathbf{x}_{t}))+g(\Pi_{\mathcal{X}_{\rho}}(\mathbf{x}_{t}))$
$\leq G_{g}\mathrm{dist}(\mathbf{x}_{t},\mathcal{X}_{\rho})-\rho$
$\leq G_{g}\gamma^{(t-1)/2}\mathrm{dist}(\mathbf{x}_{1},\mathcal{X}_{\rho})+\frac{\eta G_{g}G_{f}}{1-\sqrt{\gamma}}-\rho.$
This matches the constraint violation guarantee in Theorem 1.